OpenStudy (anonymous):

radius of a sphere is 10mm.. what is the circumference of great circle? surface of the sphere? volume of the sphere? leave answers in the terms of pi

OpenStudy (anonymous):

The formulas are as follows 1. Circumference of great circle = 2(pi)r 2. Surface of Sphere = 4(pi)r^2 3. Volume of the sphere = 4/3(pi)r^3 where r is the radius of the sphere.

OpenStudy (anonymous):

so the great circle would be =20pi surface= 400pi volume=4188.79

OpenStudy (anonymous):

Well, you have to express the units too. the circumference of great circle is 20pi mm. The surface area is 400pi mm^2 and the volume is 4000pi/3 mm^3. Always remember, a number has no significance unless you have its unit of measure next to it.

OpenStudy (anonymous):

what if you had the great circle of 36pi in, and you had to find the radius, surface, volume?

OpenStudy (anonymous):

Have you done equations? In this case, 2(pi)r = 36(pi). You'll have to find r, and then plug it into the other formulas I gave you.

OpenStudy (anonymous):

so divide 36pi by 2pi

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

=18

OpenStudy (anonymous):

yes, so the radius = 18. Now you should be able to calculate the surface area and volume using the formulas I gave you.

OpenStudy (anonymous):

surface-1296in^2

OpenStudy (anonymous):

v=24429.025

OpenStudy (anonymous):

for some reason, you're leaving out the pi in the volume. Also, you're leaving out the units (in^3).

OpenStudy (anonymous):

i ment to put pi in it..

OpenStudy (anonymous):

but is volume right? in terms of pi

OpenStudy (anonymous):

No - it is 4/3 * pi * 18^3

OpenStudy (anonymous):

so being in the terms of pi it would be=432pi in^3

OpenStudy (anonymous):

no - it is 7776*pi in^3

OpenStudy (anonymous):

okay i did 18*18 for some reason...

OpenStudy (anonymous):

no worries - but you have to memorize the formulas. And use them in an equation. If you're given the volume, then you can easily find the radius. From the radius, you can find the surface area, the circumference etc.